OpenStudy (anonymous):

Find the absolute max and absolute min of f(x)= 2x^3-5x+1 on [-1, 2]

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

do you know how to derive?

OpenStudy (anonymous):

f'(x)=6x^2 -5, let f'(x)=0, 0=6x^2-5, solve for x, 5=6x^2,5/6=x^2 x= +-sqrt(5/6), sub this back into f(x) these are the 2 points of max and min. now just pretty much look at a graph to see which is which

OpenStudy (anonymous):

Guide me further plz. I get -4 and 2 for my max and min y values

OpenStudy (anonymous):

do you have a graphing calculator? where the graph is looks like a hill is the max, where it looks like a vally is a min. which ever point corresponds to those points are the max and min.